# Joint Graph Rewiring and Feature Denoising via Spectral Resonance

### v2

@@ Abstract
In graph learning the graph and the node features both contain noisy information about the node labels. In this paper we propose joint denoising and rewiring (JDR)—an algorithm to jointly rewire the graph and denoise the features, which improves the performance of downstream node classification graph neural nets (GNNs). JDR improves the alignment between the leading eigenspaces of graph and feature matrices. To approximately solve the associated non-convex optimization problem, we propose a heuristic that efficiently handles real-world graph datasets with multiple classes and different levels of homophily or heterophily. We theoretically justify JDR in a stylized setting and verify the effectiveness of our approach through extensive experiments on synthetic and real-world graph datasets. The results show that JDR consistently outperforms existing rewiring methods on node classification using GNNs as downstream models.
@@ 1 Introduction
Graph neural networks (GNNs) are a powerful deep learning tool for graph-structured data, with applications in physics (Mandal et al., 2022; Linkerhägner et al., 2023), chemistry (Gilmer et al., 2017), biology (Gligorijević et al., 2021) and beyond (Zhou et al., 2020). Typical tasks across disciplines include graph classification (Duvenaud et al., 2015; Xu et al., 2019), node classification (Kipf and Welling, 2017; Li et al., 2019) and link prediction (Pan et al., 2022).
Graph datasets contains two distinct types of information: the graph structure and the node features. The graph encodes interactions between entities and thus the classes or communities they belong to, similarly to the features. Since many real graphs inherently contain missing and spurious links and the features are noisy, we ask a natural question: Is there a simple way to jointly rewire the graph and denoise the features which improves performance of a downstream GNN?
@@
Indeed, recent work demonstrates that rewiring the graph by judiciously adding and removing edges may improve GNN performance by facilitating information flow. In a GNN, the graph serves not only to encode interactions but also to organize computations (Battaglia et al., 2018). So even when it correctly encodes interactions it may not be an effective computational graph and rewiring it may help. Preprocessing methods rewire the graph using its geometric and spectral properties (Topping et al., 2022; Nguyen et al., 2023; Karhadkar et al., 2023). End-to-end methods (Giraldo et al., 2023; Gutteridge et al., 2023; Qian et al., 2024) dynamically rewire the graph during training leveraging both the graph and the node features. However, unlike preprocessing methods, they do not output an improved graph which restricts their interpretability and reusability. Our focus is on the preprocessing methods.
@@
Figure 1: Schematic overview of joint denoising and rewiring (JDR). In this example, we consider a noisy graph as it occurs in many different real-world scenarios, in the sense that it contains edges between and within classes and its node features are not fully aligned with the labels. The graph’s adjacency matrix $\bm{A}$ and binary node features $\bm{X}$ can be decomposed via spectral decomposition and SVD. The rewiring of $\bm{A}$ is performed by combining the information of its own eigenvectors $\bm{V}$ and the singular vectors $\bm{U}$ from $\bm{X}$. The same applies vice versa for denoising, and both are performed iteratively $K$ times. We synthesise the rewired graph $\tilde{\bm{A}}$ and the denoised features $\tilde{\bm{X}}$ by multiplying back with the final $\bm{V}_{(K)}$ and $\bm{U}_{(K)}$. To get specific properties like sparsity or binarity we can perform an Update step, e.g. by thresholding (as done here). The resulting denoised and rewired graph is displayed on the right. Its structure now better represents the communities and the first entry of the features indicates the class assignment.
@@
In this work, we propose a new rewiring scheme that also uses node features (which contain information about node labels) to produce an enhanced graph. That both the graph and the features are correlated with labels is implicit in high-quality stylized models of graphs with features, including community models such as the contextual stochastic block model (cSBM) (Deshpande et al., 2018) and neighborhood graphs on points from low-dimensional manifolds. This motivates various spectral clustering and non-linear dimensionality reduction methods (Shi and Malik, 2000; Ng et al., 2001). In the cSBM, seminal theoretical work shows that jointly leveraging the graph (stochastic block model (SBM)) and the features (a Gaussian mixture model (GMM)) improves over unsupervised clustering using either piece of information alone. However, the associated efficient inference algorithms based on belief propagation (Deshpande et al., 2018; Duranthon and Zdeborová, 2023) rely on perfect knowledge of the distribution of the cSBM and cannot be applied to arbitrary real-world data.
@@ 2.1 Preliminaries
We let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $|\mathcal{V}|=N$ nodes and an adjacency matrix $\bm{A}$. To each node we associate an $F$-dimensional feature vector and collect these vectors in the rows of matrix $\bm{X}\in\mathbb{R}^{N\times F}$. We make extensive use of the graph and feature spectra, namely the eigendecomposition $\bm{A}=\bm{V}\bm{\Lambda}\bm{V}^{T}$ and the SVD $\bm{X}=\bm{U}\bm{\Sigma}\bm{W}^{T}$, with eigen- and singular values ordered from largest to smallest. (As discussed below, in heterophilic graphs we order the eigenvalues of $\bm{A}$ according to their absolute value.) The graph Laplacian is $\bm{L}=\bm{D}-\bm{A}$, where $\bm{D}$ is the diagonal node degree matrix. For $k>2$ node classes, we use one-hot labels $\bm{y}\in\{0,1\}^{N\times k}$. We write $[L]$ for the set $\{1,2,\ldots,L\}$. In the balanced two-class case, we consider nodes to be ordered so that the first half has label $\bm{y}_{i}=-1$ and the second half $\bm{y}_{i}=1$. In semi-supervised node classification, the task is to label the nodes based on the graph ($\bm{A}$ and $\bm{X}$) and a subset of the labels $\bm{y}$. Homophilic graphs are those where nodes are more likely to connect with nodes with similar features or labels (e.g., friendship networks (McPherson et al., 2001)); heterophilic graphs are those where nodes more likely to connect with dissimilar nodes (e.g., protein interaction networks (Zhu et al., 2020)).
@@ 2.2 Motivation via the contextual stochastic block model
Contextual Stochastic Block Model. CSBMs (Deshpande et al., 2018) extend SBMs (Abbe, 2018), a community graph model, by high-dimensional node features. They have become a key generative model for studying GNNs (Baranwal et al., 2021; Wu et al., 2023; Kothapalli et al., 2023); for further pointers see Appendix A.2. We use cSBMs to build intuition about the graph rewiring and denoising problem. In a balanced 2-class SBM, the nodes are divided into two equal-sized communities with node labels $\bm{y}_{i}\in\left\{\pm 1\right\}$ [^2] This can be naturally extended to several classes of different size (Duranthon and Zdeborová, 2023).. Pairs of nodes connect independently at random, with probability $c_{\text{in}}/N$ inside communities and $c_{\text{out}}/N$ across communities.
In the sparse regime (Abbe, 2018), with average node degree $d=\mathcal{O}(1)$, it is common to parameterize probabilities as $c_{\text{in}}=d+\lambda\sqrt{d}$ and $c_{\text{out}}=d-\lambda\sqrt{d}$, where $|\lambda|$ can be seen as the signal-to-noise ratio (SNR) of the graph. The signal $\bm{X}_{i}\in\mathbb{R}^{F}$ at node $i$ comes from a GMM,
$\bm{X}_{i}=\sqrt{\frac{\mu}{N}}\bm{y}_{i}\bm{u}+\frac{\bm{z}_{i}}{\sqrt{F}},$ (1)
where $\bm{u}\sim\mathcal{N}\left(0,\bm{I}_{F}/F\right)$ is the randomly drawn mean and $\bm{z}_{i}\sim\mathcal{N}\left(0,\bm{I}_{F}\right)$ is i.i.d. Gaussian standard noise. We set $\gamma=\frac{N}{F}$ and, following Chien et al. (2021), parameterize the graphs generated from the cSBM using $\phi=\frac{2}{\pi}\arctan(\lambda\sqrt{\gamma}/\mu)$. For $\phi\to 1$ we get homophilic behavior; for $\phi\to-1$ we get heterophilic behavior. Close to either extreme the node features contain little information. For $\phi\to 0$ the graph is Erdős–Rényi and only the features contain information.
Denoising and Rewiring the cSBM. In the cSBM, $\bm{A}$ and $\bm{X}$ offer different noisy views on the labels. One can show that up to a scaling and a shift, the adjacency matrix is approximately ±$\pm\bm{y}\bm{y}^{T}+\bm{Z}_{ER},$ which means that it is approximately a rank-one matrix with labels in the range, corrupted with “Erdős–Rényi-like noise” $\bm{Z}_{ER}$ (Erdös and Rényi, 1959). It similarly follows directly from the definition that the feature matrix $\bm{X}$ is approximately (again up to a scaling and a shift) $\bm{y}\bm{u}^{T}+\bm{Z}_{G}$ where $\bm{Z}_{G}$ is white Gaussian noise. It thus makes sense to use the information from $\bm{X}$ to enhance $\bm{A}$ and vice versa. Deshpande et al. (2018) show that solving
@@
for some carefully chosen value of $b$ allows one to analyze detection bounds in unsupervised community detection with $k=2$. It is easy to see that in the high-SNR case ($\lambda$ and $\mu$ far away from the detection threshold), when the second leading eigenvector of $\bm{A}$ and the leading left singular vectors $\bm{X}$ approximately coincide with the labels, the optimal $\bm{v}^{*}$ is related to those vectors and aligned with the labels. This suggests the following rationale for denoising: (1) We can interpret the value of (2) as a measure of alignment. Since $\bm{v}^{*}$ corresponds to the labels, we can relate this measure to the quality of the label estimation. (2) We may leverage this alignment to rewire the graph and denoise the features. Namely, we could perturb $\bm{A}$ and $\bm{X}$ in a way that improves the alignment.
@@
Figure 2: An illustration of spectral alignment and resonance. In (2(a)) we plot $\bm{\hat{x}}=\bm{x}^{T}\bm{A}\bm{x}$ for different noise levels in $\bm{A}$ and $\bm{x}\in\{-1,1\}^{N}$, illustrated in the rows below. Without noise, $\bm{x}$ is exactly the label vector and $\bm{A}$ is block-diagonal. We apply multiplicative noise; namely, for each noise level, we flip the sign of a proportion of values, resulting in a random signal for ±$\pm 0.5$. We see that the norm of $\bm{\hat{x}}$ depends on the noise level. The maximum is achieved for zero noise when the second leading eigenvector of $\bm{A}$ and the signal $\bm{x}$ are perfectly aligned. In (2(b)), we consider a signal $\bm{\hat{x}}=\bm{A}\bm{x}$ for different noise levels in $\bm{A}$ and $\bm{x}$ on a graph with $20$ nodes; only a quarter of edges are shown to reduce clutter; the intra-class edges are grey; the inter-class edges are black. The largest norm is obtained for noise-free $\bm{A}$ and $\bm{x}$ (upper-left corner). The norm of $\hat{\bm{x}}$ and the separation of communities decrease along both noise axes. The inherent denoising capabilities of propagating $\bm{x}$ on a high-SNR graph (Ma et al., 2021b) are also visible, particularly in the first two rows to the right.
@@
Remark: The logic of this definition is that for a cSBM with high SNR and $k$ classes, the information about labels is indeed contained in the leading $L=k$ vectors of $\bm{V}$ and $\bm{U}$. The quantity $\mathrm{Alignment}_{L}(\bm{A},\bm{X})$ is the cosine of the angle between the subspaces spanned by the columns of $\bm{V}_{L}$ and $\bm{U}_{L}$. To denoise the features and rewire the graph, we seek to improve the alignment.
@@ 2.3 Joint Denoising and Rewiring Algorithm
JDR improves alignment. We can prove that JDR improves alignment under a stylized cSBM-like model. In fact, we can show something even stronger: the algorithm improves alignment with true labels. However, since we do not know the labels in practice, we focus on improving the alignment as defined in (4). To do this, we model the graph and features as rank-1 “label” matrices $\bm{A}=\lambda\bm{yy}^{T}+\bm{O}_{A}$ and $\bm{B}=\lambda\bm{yy}^{T}+\bm{O}_{B}$, which are corrupted by independent GOE noise $\bm{O}_{A}$ and $\bm{O}_{B}$ and $\left\|\bm{y}\right\|=1$. This is a common model in the analysis of cSBM (Deshpande et al., 2018). The Gaussian adjacency equivalence conjecture (Shi et al., 2024) suggests a generalization of the findings to the true binary adjacency case. One iteration of JDR with $L=1$ first interpolates between the leading eigenvectors $\bm{v}_{1}\left(\bm{A}\right)=\bm{v}_{A}$ and $\bm{v}_{1}\left(\bm{B}\right)=\bm{v}_{B}$; graph synthesis then yields $\bm{A}_{\eta}=\bm{A}+\lambda_{1}\left(\bm{A}\right)\left(-\bm{v}_{A}\bm{v}_{A}^{T}+\tilde{\bm{v}}_{A}\tilde{\bm{v}}_{A}^{T}\right)$ with $\tilde{\bm{v}}_{A}=(1-\eta)\bm{\bm{v}}_{A}+\mathrm{sign}(\langle\bm{\bm{v}}_{A},\bm{v}_{B}\rangle)\eta\bm{v}_{B}$.
@@
Let $\lambda>\sqrt{2}$. There exists an $\eta^{0}\in(0,1)$ such that for all $\eta\in(0,\eta^{0})$, we have
$\mathbb{P}\big{[}\left\langle\bm{v}_{1}\left(\bm{A}\right),\bm{y}\right\rangle^{2}<\langle\bm{v}_{1}(\bm{A}_{\eta}),\bm{y}\rangle^{2}\big{]}\to 1\text{ as }N\to\infty.$
In words, interpolation improves alignment of the largest eigenvector with the labels $\bm{y}$ for sufficiently large graphs. By applying the same argument to $\bm{B}$, we obtain that our method indeed improves alignment and multiple iterations approximate solutions to (4). The proof can be found in Appendix A.3.

### v3

@@ Abstract
When learning from graph data, the graph and the node features both give noisy information about the node labels. In this paper we propose an algorithm to jointly denoise the features and rewire the graph (JDR), which improves the performance of downstream node classification graph neural nets (GNNs). JDR works by aligning the leading spectral spaces of graph and feature matrices. It approximately solves the associated non-convex optimization problem in a way that handles graphs with multiple classes and different levels of homophily or heterophily. We theoretically justify JDR in a stylized setting and show that it consistently outperforms existing rewiring methods on a wide range of synthetic and real-world node classification tasks.
@@ 1 Introduction
Graph neural networks (GNNs) are a powerful deep learning tool for graph-structured data, with applications in physics [Mandal et al., 2022, Linkerhägner et al., 2023], chemistry [Gilmer et al., 2017], biology [Gligorijević et al., 2021] and beyond [Zhou et al., 2020]. Typical tasks across disciplines include graph classification [Duvenaud et al., 2015, Xu et al., 2019], node classification [Kipf and Welling, 2017, Li et al., 2019] and link prediction [Pan et al., 2022].
Graph datasets contain two distinct types of information: the graph structure and the node features. The graph encodes interactions between entities and thus the classes or communities they belong to, similarly to the features. Recent work demonstrates that rewiring the graph by judiciously adding and removing edges may improve downstream GNN performance. That work argues that in a GNN, the graph serves not only to encode interactions but also to organize message passing computations [Battaglia et al., 2018]. Even when it correctly encodes interactions it may not be an effective computational graph—rewiring it may then facilitate information flow.
@@
Graph rewiring methods can be categorized into preprocessing and end-to-end. Preprocessing methods rewire the graph by relating its geometric and spectral properties to information flow [Topping et al., 2022, Nguyen et al., 2023, Karhadkar et al., 2023]. End-to-end methods [Giraldo et al., 2023, Gutteridge et al., 2023, Qian et al., 2024] dynamically rewire the graph during training, leveraging both the graph and the node features. Unlike preprocessing methods, they do not output an improved graph which restricts their interpretability and reusability. Our focus is on the preprocessing methods.
There are thus two mechanisms that hurt performance of GNNs: (1) real-world graphs and features are noisy (the graph has spurious and missing links), and (2) geometric properties of the graph impede message passing. In this paper we focus on (1) and ask a natural question: can simple, joint feature and graph denoising improve performance of a downstream GNN?
@@
Figure 1: Schematic overview of joint denoising and rewiring (JDR). In this example, we consider a noisy graph as it occurs in many different real-world scenarios, in the sense that it contains edges between and within classes and its node features are not fully aligned with the labels. The graph’s adjacency matrix $\bm{A}$ and binary node features $\bm{X}$ are decomposed via spectral decomposition and SVD. The rewiring of $\bm{A}$ is performed by combining the information of its own eigenvectors $\bm{V}$ and the singular vectors $\bm{U}$ from $\bm{X}$. The same applies vice versa for denoising, and both are performed iteratively $K$ times. We synthesize the rewired graph $\tilde{\bm{A}}$ and the denoised features $\tilde{\bm{X}}$ by multiplying back with the final $\bm{V}_{(K)}$ and $\bm{U}_{(K)}$. To get specific properties like sparsity or binarity we can perform an Update step, e.g. by thresholding (as done here). The resulting denoised and rewired graph is displayed on the right. Its structure now better represents the communities and the first entry of the features indicates the class assignment.
@@
We propose a new rewiring scheme that also uses node features to produce an enhanced graph. We leverage the fact that both the graph and the features are correlated with the labels. This is explicit in high-quality stylized models of graphs with features, including community models such as the contextual stochastic block model (cSBM) [Deshpande et al., 2018] and neighborhood graphs on points from low-dimensional manifolds. This fact motivates various spectral clustering and non-linear dimensionality reduction methods [Shi and Malik, 2000, Ng et al., 2001]. In the cSBM, seminal theoretical work shows that jointly leveraging the graph (stochastic block model (SBM)) and the features (a Gaussian mixture model (GMM)) improves over unsupervised clustering using either piece of information alone. However, the associated efficient inference algorithms based on belief propagation [Deshpande et al., 2018, Duranthon and Zdeborová, 2023] rely on perfect knowledge of the distribution of the cSBM and cannot be applied to arbitrary real-world data.
@@ 2.1 Preliminaries
We let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $|\mathcal{V}|=N$ nodes and an adjacency matrix $\bm{A}$. To each node we associate an $F$-dimensional feature vector and collect these vectors in the rows of matrix $\bm{X}\in\mathbb{R}^{N\times F}$. We make extensive use of the graph and feature spectra, namely the eigendecomposition $\bm{A}=\bm{V}\bm{\Lambda}\bm{V}^{T}$ and the SVD $\bm{X}=\bm{U}\bm{\Sigma}\bm{W}^{T}$, with eigen- and singular values ordered from largest to smallest. (As discussed below, in heterophilic graphs we order the eigenvalues of $\bm{A}$ according to their absolute value.) The graph Laplacian is $\bm{L}=\bm{D}-\bm{A}$, where $\bm{D}$ is the diagonal node degree matrix. For $k>2$ node classes, we use one-hot labels $\bm{y}\in\{0,1\}^{N\times k}$. We write $[L]$ for the set $\{1,2,\ldots,L\}$. In the balanced two-class case, we consider nodes to be ordered so that the first half has label $\bm{y}_{i}=-1$ and the second half $\bm{y}_{i}=1$. In semi-supervised node classification, the task is to label the nodes based on the graph ($\bm{A}$ and $\bm{X}$) and a subset of the labels $\bm{y}$. Homophilic graphs are those where nodes are more likely to connect with nodes with similar features or labels (e.g., friendship networks [McPherson et al., 2001]); heterophilic graphs are those where nodes more likely to connect with dissimilar nodes (e.g., protein interaction networks [Zhu et al., 2020]).
@@ 2.2 Motivation via the contextual stochastic block model
Contextual Stochastic Block Model. CSBMs [Deshpande et al., 2018] extend SBMs [Abbe, 2018], a community graph model, by high-dimensional node features. They have become a key generative model for studying GNNs [Baranwal et al., 2021, Wu et al., 2023, Kothapalli et al., 2023]; for further pointers see Appendix A.2. We use cSBMs to build intuition about the graph rewiring and denoising problem. In a balanced 2-class SBM, the nodes are divided into two equal-sized communities with node labels $\bm{y}_{i}\in\left\{\pm 1\right\}$. Pairs of nodes connect independently at random, with probability $c_{\text{in}}/N$ inside communities and $c_{\text{out}}/N$ across communities.
In the sparse regime [Abbe, 2018], with average node degree $d=\mathcal{O}(1)$, it is common to parameterize probabilities as $c_{\text{in}}=d+\lambda\sqrt{d}$ and $c_{\text{out}}=d-\lambda\sqrt{d}$, where $|\lambda|$ can be seen as the signal-to-noise ratio (SNR) of the graph. The signal $\bm{X}_{i}\in\mathbb{R}^{F}$ at node $i$ comes from a GMM,
$\bm{X}_{i}=\sqrt{\frac{\mu}{N}}\bm{y}_{i}\bm{\xi}+\frac{\bm{z}_{i}}{\sqrt{F}},$ (1)
where $\bm{\xi}\sim\mathcal{N}\left(0,\bm{I}_{F}/F\right)$ is the randomly drawn mean and $\bm{z}_{i}\sim\mathcal{N}\left(0,\bm{I}_{F}\right)$ is i.i.d. Gaussian standard noise. We set $\gamma=\frac{N}{F}$ and, following Chien et al. [2021], parameterize the graphs generated from the cSBM using $\phi=\frac{2}{\pi}\arctan(\lambda\sqrt{\gamma}/\mu)$. For $\phi\to 1$ we get homophilic behavior; for $\phi\to-1$ we get heterophilic behavior. Close to either extreme the node features contain little information. For $\phi\to 0$ the graph is Erdős–Rényi and only the features contain information.
Denoising and Rewiring the cSBM. In the cSBM, $\bm{A}$ and $\bm{X}$ offer different noisy views on the labels. One can show that up to a scaling and a shift, the adjacency matrix is approximately ±$\pm\bm{y}\bm{y}^{T}+\bm{Z}_{ER},$ which means that it is approximately a rank-one matrix with labels in the range, corrupted with “Erdős–Rényi-like noise” $\bm{Z}_{ER}$ [Erdös and Rényi, 1959]. Another way to see this is to note that $\mathbb{E}\bm{A}=\frac{1}{2N}(c_{\text{in}}+c_{\text{out}})\bm{1}\bm{1}^{T}+\frac{1}{2N}(c_{\text{in}}-c_{\text{out}})\bm{y}\bm{y}^{T}$ (from the definition of the SBM). Since $\bm{A}$ is close to $\mathbb{E}\bm{A}$ at high SNR, the eigenvectors contain information about the labels. It similarly follows directly from the definition that the feature matrix $\bm{X}$ is (up to a scaling) $\bm{y}\bm{u}^{T}+\bm{Z}_{G}$ where $\bm{Z}_{G}$ is white Gaussian noise. It thus makes sense to use the information from $\bm{X}$ to enhance $\bm{A}$ and vice versa. Deshpande et al. [2018] show that analyzing the following optimization problem:
@@
for some carefully chosen value of $b$ allows one to characterize detection bounds in unsupervised community detection with $k=2$. It is clear from the above reasoning that in the high-SNR regime ($\lambda$ and $\mu$ far away from the detection threshold), the second leading eigenvector of $\bm{A}$ and the leading left singular vector of $\bm{X}$ approximately coincide with the labels. The optimal $\bm{v}^{*}$ is related to those vectors and aligned with the labels, since the quadratic and the bilinear form in (2) are individually maximized by the mentioned vectors. The maximizer of the linear combination of both terms therefore combines the spectral information from both matrices—the graph and the features. This suggests the following rationale for denoising: (1) We can interpret the value of (2) as a measure of alignment. Since $\bm{v}^{*}$ corresponds to the labels, we can relate this measure to the quality of the label estimation. (2) We may leverage this alignment to rewire the graph and denoise the features. Namely, we could perturb $\bm{A}$ and $\bm{X}$ in a way that improves the alignment.
@@
Figure 2: An illustration of spectral alignment and resonance. In (2(a)) we plot $r=\bm{x}^{T}\bm{A}\bm{x}$ for different noise levels in $\bm{A}$ and $\bm{x}\in\{-1,1\}^{N}$, illustrated in the rows below. Without noise, $\bm{x}$ is exactly the label vector and $\bm{A}$ is block-diagonal. We apply multiplicative noise; namely, for each noise level, we flip the sign of a proportion of values, resulting in a random signal for ±$\pm 0.5$. We see that the value of $r$ depends on the noise level. The maximum is achieved for zero noise when the second leading eigenvector of $\bm{A}$ and the signal $\bm{x}$ are perfectly aligned. In (2(b)), we consider a signal $\bm{\hat{x}}=\bm{A}\bm{x}$ for different noise levels in $\bm{A}$ and $\bm{x}$ on a graph with $20$ nodes; only a quarter of edges are shown to reduce clutter; the intra-class edges are grey; the inter-class edges are black. The largest norm is obtained for noise-free $\bm{A}$ and $\bm{x}$ (upper-left corner). The norm of $\hat{\bm{x}}$ and the separation of communities decrease along both noise axes. The inherent denoising capabilities of propagating $\bm{x}$ on a high-SNR graph [Ma et al., 2021b] are also visible, particularly in the first two rows to the right.
@@
Remark: The logic of this definition is that for a cSBM with high SNR and $k$ classes, the information about labels is indeed contained in the leading $L=k$ vectors of $\bm{V}$ and $\bm{U}$. This follows directly by generalizing the formulation in (2) to multiple classes and thus multiple eigenvectors [Decelle et al., 2011, Lesieur et al., 2017]. The quantity $\mathrm{Alignment}_{L}(\bm{A},\bm{X})$ is the cosine of the angle between the subspaces spanned by the columns of $\bm{V}_{L}$ and $\bm{U}_{L}$. To denoise the features and rewire the graph, we seek to improve the alignment.
@@ 2.3 Joint Denoising and Rewiring Algorithm
JDR Improves Alignment. We now show that JDR improves alignment, as defined in (4), under a stylized cSBM-like model. In fact, we show a stronger result: that the algorithm improves alignment with the true labels. Appealing to universality arguments [Hu and Lu, 2023], we study a model with a spiked Gaussian matrix $\bm{A}^{c}=\frac{\lambda}{N}\bm{yy}^{T}+\frac{1}{\sqrt{N}}\bm{O}_{A}$ where $\bm{O}_{A}$ is GOE noise instead of the binary matrix $\bm{A}$, and features $\bm{X}=\sqrt{\frac{\mu}{N}}\bm{y\xi}^{T}+\frac{1}{\sqrt{F}}\bm{O}_{X}$ as defined in 1. In the cSBM context, a recent conjecture with strong empirical support states that replacing the binary $\bm{A}$ by $\bm{A}^{c}$ leads to the same behavior in downstream tasks such as community detection [Deshpande et al., 2018, Lu and Sen, 2023] and node classification [Shi et al., 2024]. An iteration of JDR with $L=1$ applied to this model, first interpolates between the leading eigenvectors $\bm{v}_{1}\left(\bm{A}^{c}\right)=\bm{v}_{A}$ and leading left singular vector $\bm{u}_{1}\left(\bm{X}\right)=\bm{u}_{X}$. Graph and feature synthesis then yields $\bm{A}^{c}_{\eta_{A}}=\bm{A}^{c}+\lambda_{1}\left(\bm{A}^{c}\right)\left(-\bm{v}_{A}\bm{v}_{A}^{T}+\tilde{\bm{v}}_{A}\tilde{\bm{v}}_{A}^{T}\right)$ and $\bm{X}_{\eta_{X}}=\bm{X}+\sigma_{1}\left(\bm{X}\right)\left(-\bm{u}_{X}\bm{w}_{X}^{T}+\tilde{\bm{u}}_{X}\bm{w}_{X}^{T}\right)$. Here $\tilde{\bm{v}}_{A}=(1-\eta_{A})\bm{\bm{v}}_{A}+\mathrm{sign}(\langle\bm{\bm{v}}_{A},\bm{u}_{X}\rangle)\eta_{A}\bm{v}_{X}$ and $\tilde{\bm{u}}_{X}=(1-\eta_{X})\bm{u}_{X}+\mathrm{sign}(\langle\bm{v}_{A},\bm{u}_{X}\rangle)\eta_{X}\bm{v}_{A}$, where $\bm{w}_{1}\left(\bm{X}\right)=\bm{w}_{X}$ is the leading right singular vector of $\bm{X}$. Denoting $\tilde{\bm{y}}=\bm{y}/\sqrt{N}$, we have
@@
Let $\lambda>1$ and $\mu>\sqrt{\gamma}$ with $\gamma=N/F$. There exist $\eta_{A}^{0},\eta_{B}^{0}\in(0,1)$ such that for all $\eta_{A}\in(0,\eta_{A}^{0})$ and $\eta_{B}\in(0,\eta_{B}^{0})$, when $N\to\infty$, we have
$\left\langle\bm{v}_{1}({\bm{A}^{c}_{\eta_{A}}}),\tilde{\bm{y}}\right\rangle^{2}\overset{a.s}{>}\left\langle\bm{v}_{1}({\bm{A}^{c}}),\tilde{\bm{y}}\right\rangle^{2}\quad\text{and}\quad\left\langle\bm{u}_{1}({\bm{X}_{\eta_{X}}}),\tilde{\bm{y}}\right\rangle^{2}\overset{a.s}{>}\left\langle\bm{u}_{1}({\bm{X}}),\tilde{\bm{y}}\right\rangle^{2}.$
In words, interpolation improves alignment of the largest eigenvector with the labels $\bm{y}$ for sufficiently large graphs. The proof, based on the BBP transition in the spiked covariance model [Baik et al., 2005] and the fluctuation of the leading eigenvector, can be found in Appendix A.3. It seems challenging but quite possible to extend this argument to a binary $\bm{A}$. One would then interpolate between $\bm{u}_{x}$ and $\bm{A}$’s second leading eigenvector $\bm{v}_{2}(\bm{A})$, which has similar properties to $\bm{v}_{1}(\bm{A^{c}})$, especially in a dense graph regime [Nadakuditi and Newman, 2012].